OpenStudy (anonymous):

Reduce the following rational expression to lowest terms, if possible. 8x+40/8x^2+72x+160

OpenStudy (anonymous):

\[\frac{ 8x+40 }{ 8x^2+72+160 }\] \[\frac{ 8(x+5 }{ 8(x^2+9x+20) }\] \[8(x+5)(x+4)\]

OpenStudy (anonymous):

Then both 8(x+5) cancels out & I'm left with: \[\frac{ 1 }{(x+4) }\]

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@ganeshie8 , is this correct?

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

Really?! omg! yaaaay! lol